OpenStudy (anonymous):

If a right triangle has legs of lengths a and b, then the length of its hypotenuse can be found using the expression √a^2 + b^2. Find the length of the hypotenuse of a right triangle whose legs measure 11cm and 14cm. Round your answer to the nearest tenth

OpenStudy (anonymous):

you can also use a^2+b^2=c^2 11^2+14+2=c^2 121+196=c^2 317=c^2 √317=c 17.8=c

OpenStudy (anonymous):

=17.8